OpenStudy (anonymous):

Fill in the missing term so that the quadratic equation has a graph that opens up, with a vertex of (–2, – 16), and x intercepts at x = –6 and x = 2. (Do not include the negative sign in your answer.) y = x2 + 4x − ___ Help me on this one it's hard

OpenStudy (anonymous):

well the answer is y=x^2+4x+12

OpenStudy (anonymous):

Well. First of all, Vertex is (-b/2a, -D/4a) Let the constant be some C. Then on satisfying -2 in the equation, => y = -2*-2 + -2*4 + c We know that at this point y = -16. => -16 = 4 - 8 + c = > c = -12

OpenStudy (anonymous):

Rest of the information is simply for you to confirm your answer. Which is c = 12 Equation being, x^2 + 4x - 12.

OpenStudy (anonymous):

cuz we have given the vertex ie (-2,-16) wichich is also equal to (-b/2a, 4ac-b^2/4a) and by comparing the equation with ax^2+bx+c we'll get the answer

OpenStudy (anonymous):

well there won't be '-" sign cause there is already one siddhantsharan

OpenStudy (anonymous):

Yeah. But I assumed the equation as ax^2 + bx +c. So I get c = -12. => Equation has a The constant term as -12. Though as I specified above too, c = 12. ONLY because - sign is provided. Your equation in the first post is incorrect. Check, it doesnt satisfy the roots.

OpenStudy (anonymous):

well i guess you were right siddhantsharan